OpenStudy (anonymous):

Simplify 13 + 2 - 5 - 15 -25 -5 5

OpenStudy (anonymous):

pemdas 13 + 2 - 5 - 15 15 - 5 - 15 10 - 15 answer: -5

OpenStudy (anonymous):

-5 Simply use order of operations you only have addition and subtraction so work left to right

OpenStudy (anonymous):

13+2=15 -5+-15=-20 so 15+-20 = -5 answer:-5